OVHcloud Network Status

Current status
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#8007 — Link Roubaix<>London 05-06/02
Scheduled Maintenance Report for Network & Infrastructure
We are accomplishing the final step to move the optical link Roubaix<>London which is switching on the technology 100G coherent. This step implies switching to new cards BMM2C to extremity of the link as well as replacing all amplification cards on the intermediate sites (2 in FR, 2 in UK) which are passing in RAMAN. These works will allow us to touch directly Amsterdam from Roubaix without regenerating signal which will allow securing the circuits Roubaix<>Amsterdam (th circuits will switch on a path via London in case of a cable cutoff between Roubaix and Amsterdam).
the intervention will take a long time due to fixing the RAMAN amplis, We are expecting about 4 to 5 hours of cutoff with a booting at 00:00 CET.


Date: 2013-02-06 07:21:35 UTC
Traffic restored

Date: 2013-02-06 07:19:18 UTC
We have reactivated the circuits Roubaix <> London. This link works now by 100G coherent. We still need to work on London to fix the problems on the last meters of the fiber, but this operation can be done when the protection by Amsterdam will be operational and seamless.

Date: 2013-02-06 07:17:06 UTC
We will restore 10G circuits.

Date: 2013-02-06 07:13:48 UTC
We were not able to fix the issue on the fiber London's side but this is not a problem for remounting the link back.
We are largely in the nails.
We will intervene another night to work with Interoute on this part.

We will finish all the settings on the last stretch Wittes> Roubaix then reactivate 10G circuits between Roubaix and London.

Date: 2013-02-06 07:08:05 UTC
We will not be able to find the problem on RX London tonight. It is already> 6H and it is time to complete operations.
We'll restore the circuits as soon as London/Roubaix is completed. Then we will schedule an intervention another night in order to improve RX span between different DC at London where we have no access to tonight.

Date: 2013-02-06 06:39:33 UTC
We still have RX problem on London's side . We are starting the configuration script RAMAN in the direction London> Roubaix.
Two engineer went to Globalswitch en// trying to fix the problem there.

Date: 2013-02-06 06:36:12 UTC
The script should run with direction Roubaix> London then the direction London> Roubaix which will take 1 hour in each direction.

The direction Roubaix> London just ended but the value of loss point measured by the script in London is very high which means that there is an issue about the wiring in the first meters in London. The team is currently on site to check the connections.

Date: 2013-02-06 06:33:31 UTC
All cards are in place and patchings is ok. We can already say at this stage that the optical perfs on the link are good and the link should be able to link up smoothly.
Spans setting scripts on RAMAN are however running and we must keep the circuits cut until the operation is done.

Date: 2013-02-06 06:29:27 UTC
Migration started.

Date: 2013-02-06 06:29:01 UTC
During 5H we will have lost packets between our DC and RBX UK, U.S. and CA.

Date: 2013-02-06 06:28:23 UTC
Teams are on the pitch at each site (the intervention tonight involves fifteen persons on the path of the link). Early intervention is imminent.
Posted Feb 04, 2013 - 23:45 UTC